Sr. financial analyst, Mindy Jo, has forecasted the free cash flows of Southern Sweets to grow by 14.00% next year, 8.00% in the following year and 7.00% in the third year before leveling off to a constant growth rate of 3.00% annually. The firm’s most recently reported free cash flow was $126,000.00. The firm has $6,500.00 in marketable securities, $90,000.00 in long-term debt, $15,000.00 in preferred stock, and $52,000.00 in book common equity. Southern Sweets has 40,000.00 common shares outstanding and a cost of capital of 11.00%.

What is the value of operations for Southern Sweets?
Answer #1

Value of operations =Present Value of Free cash flows

Present Value of cash flow=(cash flow)/((1+R)^N))

R=Cost of capital =11%=0.11

N=Year of cash Flow

Recent Free Cash Flow=F0=$126,000

Free Cash Flow of year1=F1=126000*(1+0.14)=$143,640

Free Cash Flow of year2=F2=$143,640*(1+0.08)=$155,131

Free Cash Flow of year3=F3=$155,131*(1+0.07)=$165,990

Free Cash Flow of year4=F4=$165,990*(1+0.03)=$170,970

Horizon Value of Free Cash Flows from Year 4 to Infinity at 3% constant Growth Rate (g) , in Year 3

HF3=F4/(R-g)=170970/(0.11-0.03)=170970/0.08=$2,137,126

PV1=Present Value of Horizon Value in year 3=HF3/((1+R)^3)=2137126/(1.11^3)=$1,562,648

PV2=Present Value of Free Cash Flow of Year 3=F3/((1+R)^3)=165990/(1.11^3)=$121,371

PV3=Present Value of Free Cash Flow of Year 2=F2/((1+R)^2)=155131/(1.11^2)=$125,908

PV4=Present Value of Free Cash Flow of Year 1=F1/((1+R)^1)=143640/1.11=$129,405

Value of Operations =PV1+PV2+PV3+PV4=$1,939,332

Value of operations $1,939,332